14. Tobramycin 5 mg IV is ordered for an infant who weighs 5lbs. The recommended dose is 2 to 2.5 mg/kg q 8hrs.

a. Is this a therapeutic dose for the child? _______
Answer #1

ordered dose=5 mg

weight= 5 lbs=2.267 kg ( as 1 lbs=0.4535 kg )

Recommended range of drug=2 - 2.5 mg /kg

Dose for child=weight \times recommended range of drug

=2.267 \times 2 OR 2.267 \times 2.5

=4.534 OR 5.6675 mg

safe range of drug =4.5-6 mg / kg ( approximately)

therefore this ordered dose of 5 mg is therapeutic dose for child.